Journalist Report – May 4th

Crew 235 Journalist Report 04May2021
Author: Krysta Darby, Crew Journalist
Sol 2: Rock Sampling and Mission Patch Design

The day started at approximately 7:15am Earth’s MT when Thomas got up and started the coffee. Members of the crew got up one by one, eager to look out onto the Martian landscape. Atila filled the water tank, but recommended going forward it be filled at night in case there are any unforeseen issues overnight. A rodentia unfamilia was captured at an unknown time and was released out of the airlock by Randall. Director Shannon continued to work on the new generator, and the crew went into low power mode. While Krysta fixed a breakfast of spinach scrambled eggs and spam, the crew assembled around the table for morning coffee. Conversation over breakfast ranged from the day’s tasks to human rights issues in education. So far this crew has not shied away from digging deep into controversial topics. Atila and Randall led cleaning efforts in the kitchen as breakfast and a heated debate came to a conclusion.

After breakfast, Atila briefly reviewed the map of local terrain with the crew.

Today’s goal was for both teams to retrieve rock samples from Kissing Camel.

At 9:00, the crew traveled to the science dome to review the types of samples that could typically be found in the surrounding area with our resident expert, Shannon. For the purposes of this expedition, we will be referring to petrified wood as sulfur, blueberries as hematite, and lava rock as breccia. Temperature of the incubator was found to be 56 degrees Celsius. Atila adjusted to bring the temperature down to 25 degrees.

At 9:45, the crew headed back to the hub to prepare for the first team’s expedition. Jen, Randall, Jeff, and Allison stayed behind and prepared a lunch of macaroni and cheese casserole, while Atila, Krysta, Thomas, and Jennifer left the hub to collect samples from Kissing Camel. Krysta and Jennifer rode ATV’s and Atila and Thomas took the rover. They successfully collected several samples of sulfur along with some unique samples which will be analyzed. One fatality was reported due to a removed glove. Crew members are still acclimating to wearing suits for extended periods of time. A lack of maneuverability and the weight of the suit has been a challenge, but the adrenaline from inhabiting Mars fuels the team making the weight and discomfort of a space suit a minor inconvenience.

The first team returned to the hub at approximately 11:50am, and assisted the second team as they prepared for their expedition. First team enjoyed the macaroni and cheese casserole heartily and cleaned up the aftermath. The second crew reached Kissing Camel with no issues and collected samples, taking careful note of the context in which each sample was found. There was a second fatal mistake while the commander was adjusting her helmet. The crew is learning fast. Second team returned at about 1:30pm. They then had to sanitize the helmets and return the suits.

After comparing rock finds, the crew collaborated on patch design. It was noted that patch designs typically have the mission number/name/designation, and the design should represent the purpose of the mission. Last names of crew members usually border the patch, and flags of nationalities are also incorporated. After looking at examples and some brainstorming, the crew decided to incorporate symbolic representations of education, STEM, and collaboration. The patch will also incorporate a hexagon to represent the Beehive State, geometric lines representing the geodesic science dome, and the overall shape will represent the habitat. Allison has taken lead in final design as the person with the most experience in this field.

Each Crew member is currently working on their own project (reports, patch design, and telescope observations) for a few hours before the crew will come together for team building and dinner.

EVA Report – May 4th

Crew 235 EVA Report 04 May 2021

EVA # 2, EVA #3

Author: Jen Carver-Hunter

Purpose of EVA: Specimen collection

EVA 2 Start time: 10:00

EVA 2 End time: 12:00

EVA 3 Start Time: 12:00

EVA 3 End time: 1:30

Narrative: One of our crew’s goals is to collect rock samples that replicate the rocks found on Mars. For today’s EVA, our targeted specimen was petrified wood, which we are using as an analogous specimen for the hematite found on Mars. Because Kissing Camel Ridge has an excellent source of petrified wood, we chose that location for our EVA.

The crew left the hab in vehicles and followed the road south to Kissing Camel Ridge. Crew walked to the southwest side of Cow Dung Road and followed the wash to the west along the ridge. Crew continued along the wash to the source of the petrified wood specimens in the wash. The crew on EVA 2 found a location that they thought might be the source Dr Rupert mentioned. It looked like someone had piled up mounds of petrified wood.

Destination: Kissing Camel Ridge

Coordinates (use UTM NAD27 CONUS): 518250, 4249250

EVA 2 Participants: Atila, Krysta, Jennifer, and Thomas

EVA 3 Participants: Jen, Allison, Jeff, and Randall

Road(s) and routes per MDRS Map: Entrance Road to Cow Dung Road

Mode of travel: Pathfinder Rover and 2 ATVs

Operations Report – May 4th

Crew 235 Operations Report 4-MAY-2021

SOL: 2

Name of person filing report: Atila Meszaros

Non-nominal systems: Musk Observatory and Power System

Notes on non-nominal systems: Nothing to report

Generator: Working nominally. The battery is charging through the battery charger connected from the Science Dome.

Hours run: 0

From what time last night: the generator didn’t run

To what time this morning: the generator didn’t run

List any additional daytime hours when the generator was run: N/A

Solar— SOC 90% at 7:01 pm (Before generator is run at night)

Notes on power system: This morning the charge controllers were down, breakers jumped. We have been having energy through the day, but we went into power save mode during the morning. Shannon fixed the system in the morning and the SOC jumped again to 100% this afternoon. It has been steady since then.

Diesel Reading – 47%

Station Propane Reading – 80%

Outpost Propane Reading – 1. 72% 2. 80%

Water (loft tank): 25 gallons

Water Meter: 150994.0 units

Water (static tank): 400 gallons

Static to Loft Pump used – yes

Water in GreenHab: 0 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: Yes

Sojourner rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Spirit rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Opportunity rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Curiosity rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Perseverance rover used: Yes

Hours: 208.

Beginning charge: 100%

Ending charge: 99%

Currently charging: Yes

Notes on rovers: None

ATV’s Used: Yes

Reason for use: EVA#2 and EVA#3

Oil Added? No

ATV Fuel Used: N/A

# Hours the ATVs were used today: 0

Notes on ATVs: ATV’s will be filled with Free Ethanol Gas to be ready for tomorrow’s EVA.

HabCar used and why, where?: Nothing to report

CrewCar used and why, where?: Nothing to report

General notes and comments: Nothing to report

Summary of internet: Working nominally.

Summary of suits and radios: All radios nominal, all functional suits currently charging.

Summary of Hab operations: We identified a small leak under the kitchen sink, but it was just a pipe connection that has already been connected and secure. It looks like there’s some water activity in the toilet room. The floor is getting constantly wet every time we use it. We caught a little rodent during the morning and it was released in Kissing Camels.

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Geology Class in the morning with Shannon.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Rotation system won’t work in the Musk Observatory. Waiting for instructions from Peter.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Thank you for being with us tonight!

Sol Summary #2 – May 4th

Crew 235 Sol Summary Report 04-May-2021

Sol: 2

Summary Title: Jurassic Hunting

Author’s name: Atila Meszaros

Mission Status: Crew Biologist

Sol Activity Summary:

Finally on Mars! It’s been quite a while for some of us: our Commander’s last mission and mine was the last cohort for the Spaceward Bound in 2019, and after more than a year and a half, it feels wonderful to be back to Sim. Today was our first day in Sim. After a wonderful Geology lesson class from Dr. Rupert, we suited up and embarked to our destination to hunt petrified wood in the southern region of the Kissing Camels, a late Jurassic deposit. After our two successful EVAs, we did a small rock show between each other to see our bounties. We are preparing dinner and getting ready for a lovely night.

Look Ahead Plan: Our geology hunting will get more in-depth. Will visit two new sites to find a series of different rock samples. We would get into some hardcore Lego building and more teambuilding activities.

Anomalies in work: None.

Weather: Fully sunny, but with a great temperature throughout the day.

Crew Physical Status: Everyone is healthy, hydrating, and eating pretty good.

EVA: Tomorrow EVA #4 and EVA #5 to continue or geology hunting!

Reports to be filed: Operations Report, EVA Report, EVA Request, Crew Journalist

Support Requested: Nothing to report.

Copyright © The Mars Society. All rights reserved. | Main Site